# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v1

@@ 5.2 Impact of Model Settings (RQ2)
Recency of Source Material in Prompts. The models we tested were shown to be more sensitive to hallucinations when coding prompts that deal with more recent topics were used. The best performing model for this experiment was WizardCoder 33B, with a difference between the rate of recent prompts and all-time prompts of only 0.58%, with the more recent dataset resulting in a hallucination rate of 14.59% compared to 14.01% for the all-time dataset. Representing the commercial models, GPT-4 Turbo was a close second with a difference of 0.63% between recent and non-time constrained prompts. Note that a lower difference between the rates of recent and all-time prompts indicates better performance in handling questions that fall outside the model’s pre-training data. Notably, WizardCoder, a fine-tuned version of the DeepSeek models, achieved a 50% lower difference in hallucination rates during the recency experiment than the DeepSeek models after recording similar results in the primary experiment. WizardCoder uses a novel fine-tuning approach, named Evol-Instruct, that uses existing open-source data to create additional datasets that are more complex, diverse, and realistic [51]. While the primary goal of these fine-tuning methods is to increase overall coding proficiency, the results, especially achieving better performance than the GPT models in recency testing, indicate this process has a secondary benefit in reducing hallucinations that is worthy of future work. Overall, all 16 Python models we evaluated demonstrated a higher hallucination rate when being prompted about questions or packages that were popular within the past year (Figure 3). These higher rates are partly due to inherent limitations of LLMs. As noted in the OpenAI GPT-4 technical report [1], LLMs cannot update themselves with new information post-release and struggle with prompts beyond their training data cut-off date. This cut-off date is critical because models lack knowledge of events that occur after it. Although fine-tuning can enhance specific tasks, it does not generally improve the model’s overall knowledge of the world. The extensive testing and fine-tuning required for modern foundational models create a significant lag between the training data cut-off and model release, widening the gap in knowledge of recent information. The massive cost of training LLMs from scratch, indicated by the 1,400,000 GPU hours (220 years) required for the 12 CodeLlama models, makes continuously updating pre-training data impractical [71]. This cost, along with increasing model sizes and training times, poses a significant barrier to reducing package hallucinations for advanced coding prompts and packages.
Figure 3: Hallucination rates of recent vs. all-time data sets.
Effect of Temperature Settings. We varied the temperature (Section 2.1) setting for each model between the minimum and maximum allowed values and observed the change in hallucination rate (the maximum temperature for GPT series models is limited to 2, while the open-source models can be set up to 5). All models exhibited a clear increase in hallucination rate as temperature value increases, with the effect becoming severe at maximum values. The OpenAI models, as shown in Figure 4, displayed only a slight increase in hallucination rate between temperatures 0 and 1, which then increased sharply between 1 and 2. Most notably, GPT-4 resulted in a hallucination rate of (8.9%), nearly 4x lower than GPT 3.5 (31.8%) at its maximum temperature. At the highest temperature values, open-source models start to generate more hallucinated packages than valid packages. Further, open-source models such as CodeLlama, also demonstrated a tendency to reduce hallucination rate from temperature values 0 to 1.5 before beginning a rapid ascent from 1.5 to 5 (see Figure 4). Most LLMs operate at a default temperature in the range of 0.7 to 1, which by our results is slightly sub-optimal from a package hallucination perspective. Users of commercial LLMs do not have control over this parameter unless they are using the API, which incurs an additional cost beyond a monthly subscription. Consequently, open-source users have more fine-grained control of temperature values, although the proper use of the temperature setting depends on user knowledge, as well as the specific model and environment being employed.
@@
Effect of Decoding Strategies. Next, we adjusted several decoding parameters (top-$p$, top-$k$, and min-$p$ values) to reduce the chances of a low probability token being selected as a potential package, with the intuition that lower probability tokens correspond to higher probabilities of hallucination in this context. Below is a summary of the parameters and the values we modified. We evaluated each listed value in isolation, followed by a combined evaluation of the values highlighted in bold, resulting in a total of 10 tests. Note that top-$k$ and min-$p$ were only tested for DeepSeek and CodeLlama, as those values are not able to be modified through the OpenAI API.
@@
Varying the decoding values induced a slight increase (1.16% on average) in hallucination rate for all four models across all values tested. As we will expand on in RQ3, package hallucinations are often persistently repeated across many iterations. This suggests that even greedy decoding strategies, which prioritize the most probable tokens (i.e., the most probable token is always selected), can generate fictitious packages. The persistent nature of package hallucinations highlights the inherent complexity of the problem and underscores that this phenomenon is a deeply embedded in the fabric of these models.
Figure 5: (a) Frequency of the occurrence of the same hallucinated package when querying a model with the same prompt 10 times, (b) No. of models in which each unique package (valid and hallucinated) appeared, with the y-axis on a log scale, (c) No. of hallucinated packages vs. Levenshtein to the nearest valid package.
@@ 5.3 Model Behaviors (RQ3)
Frequency of Repeated Hallucinations. To determine whether hallucinations are random error or repeatable phenomena, this analysis will focus on the persistence of hallucinations within a model. We randomly sampled 500 prompts that generated package hallucinations during our initial testing and repeated those queries 10 times per prompt, then recorded how many times the original hallucination was regenerated. Our analysis reveals an unexpected dichotomy when repeatedly querying a model with the same prompt that generated a hallucination: a certain set of hallucinated packages were repeated in all 10 iterations (43%), while another set did not repeat at all across the 10 iterations (39%). This is indicated in Figure 5a, which shows notable spikes at zero repetitions and at 10 repetitions respectively for all the models. Further, 58% of the time, a hallucinated package is repeated more than once in 10 iterations, which shows that a majority of hallucinations are not simply random errors, but a repeatable phenomenon that persists across multiple iterations. This is significant because a persistent hallucination is more valuable for malicious actors looking to exploit this vulnerability and makes the hallucination attack vector a more viable threat.
Verbose Models vs. Conservative Models. LLMs operate with inherent randomness and uncertainty, yet this uncertainty is generally not communicated by the LLM to users. Rather, the results are usually presented confidently without mentioning of potential fact-conflicting statements. This non-deterministic behavior of LLMs enables novel and creative output, a desired feature for many NLP tasks but less welcome for code generation, which requires a high degree of accuracy and must adhere to rigid syntax. We define a verbose model as one that operates with higher degree of uncertainty and randomness by generating a greater number of distinct package names while a conservative model generates a lesser number of distinct packages, generally using only the most popular and well-known packages. To this end, we investigated whether verbose models correspond to higher rate of package hallucinations. Our results show a correlation between the hallucination rate and number of unique packages that were recommended during this experiment (i.e. a more verbose model was associated with a higher hallucination rate). In light of these findings, as shown in Figure 6, it is reasonable to suggest that coding models, or models performing coding tasks, should operate in a more conservative manner when generating packages to answer a coding prompt.
@@
Curiously, 3 of the 4 models (GPT 4 Turbo, GPT 3.5, and DeepSeek) proved to be highly adept at detecting their own hallucinations (see Figure 7) with detection accuracies above 75%. The precision and recall values for this test, which also averaged above 75%, are presented in Appendix F. This phenomenon implies that each model’s specific error patterns are detectable by the same mechanisms that generate them, suggesting an inherent self-regulatory capability. The indication that these models have an implicit understanding of their own generative patterns which could be leveraged for self-improvement is an important finding for developing mitigation strategies. While capable of detection, none of the models displayed a high enough detection accuracy (all less than 90% cumulative accuracy) to be relied upon as a stand-alone classification mechanism to distinguish between hallucinated vs. valid packages.
Across the board, each model displayed a better ability to identify valid vs. hallucinated packages generated by other models rather than its own, suggesting each model has certain “blind spots” inherent to their architecture. CodeLlama displays unique and interesting behavior during both the tests, as it has an overwhelming propensity to label most packages as valid, resulting in a lower accuracy for hallucinated packages (see Figure 7).
@@ 5.4 Characteristics of Hallucinations (RQ4)
Occurrence of the Same Hallucination Across Different Models. To analyze the possibility of the same hallucinated packages being generated across different LLM models, we measured how many other models also generated the same package name given a confirmed package hallucination. Figure 5b shows a clear pattern where a vast majority of distinctly generated package names were generated by only one model. In other words, specific package names were usually unique to a single model, where only the most common packages were generated by more than one model. The two populations diverge as the number of models increases, with the number of hallucinated packages decreasing nearly exponentially and the distribution of valid packages becoming more uniform. The finding that valid packages are less dependent on the specific model used for generation can be accounted to the widespread use of the most popular packages. These packages frequently appear in training data and are applicable to the most universal coding problems, causing them to be used in a significant portion of prompts.
@@
Combining the insights gained during the previously discussed persistence analysis (Figure 5a) leads to a key observation. Hallucinations are often persistent (61% are repeated within 10 iterations) within the same model but are not often repeated between models, as 81% of hallucinated packages are generated by only one model. This further reinforces the evidence that while hallucinations are a common phenomenon across various models, the exact nature of these hallucinations is generally model-specific. This suggests a ubiquitous susceptibility to hallucinations in the architecture or training process of these models, but the manifestation of such errors is influenced by each model’s unique configuration and dataset.
@@
Semantic Similarity Between Hallucinated and popular packages. In order to analyze the semantic similarity between hallucinated and popular real/valid packages, we measured the average Levenshtein distance of a package to its nearest neighbor (i.e., the closest valid package). Levenshtein distance is a measure of how many insertions, deletions, and substitutions are required for two strings to match [38]. If the distribution of Levenshtein distances is skewed heavily right, with a peak at or near 0, this would indicate that most hallucinations are very similar to valid package names. In that case, attackers could infer a hallucination target based on more traditional package confusion methods (e.g., typosquatting) rather than analyzing a large volume of model output over time to detect persistent hallucinations that could be used as vessels for malicious code. A higher distance reflects that package hallucinations are more random in nature and difficult to predict, rather than the result of minor grammatical errors. Due to the computational complexity of calculating Levenshtein distance for each hallucinated package compared to all valid Python packages, we limited our search to the most downloaded 20k Python packages according to the PyPI BigQuery dataset [66]. This approach mimics traditional package confusion methods, since typosquatting a rarely used package is not lucrative for attackers.
@@
The results of the Levenshtein distance test, as seen in Figure 5c, suggests that most package hallucinations are not simple off-by-one errors. An off-by-one error in our case is defined as a difference between the hallucinated package and its closest match of 1-2 numbers, letters, or punctuation marks, with only 13.4% (10,263 of 76,489) belonging to this category. Another 37.9% (29,025 of 76,489) of packages registered a score between 3-5, which would indicate two words with a common root word or concept that still differ significantly. Notably, 48.6% (37,207 of 76,489) of hallucinations scored 6 or greater, with 20.2% (15,457 of 76,489) of those being 10+, indicating two strings that are very different and likely do not share any common theme.
@@
The presence of such a large proportion of high Levenshtein values suggests that the majority of hallucinations are not merely trivial typographical errors but are instead substantively different from existing package names. The observed results provide further evidence that the root cause of hallucinations is likely to be more complex than minor string manipulation, pointing to deeper issues in the model’s generative processes that govern the creation of package names. The long right tail of the distribution in Figure 5c indicates a wide variety of hallucinations spread across a broad range, revealing a diversity in types of errors and reinforcing that the generation of hallucinations is a complex issue not limited to simple character substitutions, additions, and deletions.
@@
Effect of Language Confusion. Another interesting behavior observed during the main experiment is that many models have a tendency to confuse programming languages while generating package output. In other words, the model is asked to provide Python packages, but instead provides JavaScript packages. To validate the potential existence of this behavior in code generating LLMs, we obtained master lists of packages (using libraries.io [78]) from the nine most popular open-source repositories and compared our list of hallucinated package names generated during Python testing to the respective master lists of valid packages from other programming languages. Any intersection between the two lists would indicate that the LLM, while being asked for Python code, generated a package name that is not published to PyPI but is a valid package name in another programming language (i.e. the LLM generated a package for the wrong language). Overall, only JavaScript is a significant source of cross-language hallucinations, as 8.7% (6,705 of 76,489) of hallucinated Python packages are valid JavaScript packages. All other languages contributed negligible hallucinations, combining for only 0.8% (663 of 76,489) across eight other open-source repositories which includes R, Rust, Ruby, PHP, Swift and .NET (see Appendix E for complete results).
@@ 6 Mitigation
Motivated by the finding of systemic package hallucination across all the models we evaluated, we then investigated methods to mitigate the occurrence of hallucinations (addressing RQ5) by incorporating several of the most popular techniques from the literature.
@@ 6.1 Mitigation Strategies
A seemingly straightforward mitigation strategy when detecting package hallucinations might involve comparing a master list of valid packages to the model’s code output. It is imperative to understand why this approach is merely a superficial solution rather than addressing the core issue at hand. The fundamental vulnerability inherent in code generation by LLMs regarding package hallucinations is not solely that they produce non-existent packages, but that they generate package names that were not present in their training data. Our study emphasizes the detection of hallucinated packages as indicators of an underlying vulnerability. The genuine threat posed by package hallucinations emerges when such a hallucinated package is officially published and contains malicious code. The mere act of publishing does not render a previously hallucinated package benign; on the contrary, it becomes an active threat seeking to exploit a security vulnerability. Consequently, simply cross-referencing generated packages with a master list would erroneously validate a malicious package post-publication. Dedicated mitigation techniques are essential to prevent the generation of hallucinated packages altogether, thereby addressing the root cause of this latent vulnerability.
Implementing mitigation strategies specific to code generation LLMs is an unexplored topic, therefore we rely on general hallucination mitigation strategies proposed for standard NLP tasks that can be applied to code generation. These strategies can generally be grouped into two broad categories: prompt engineering and developing models [79]. Prompt engineering includes methods such as Retrieval Augmented Generation (RAG), self-refinement, and prompt tuning. RAG approaches involve enriching the original prompt with additional information gathered from an external source, such as the web or a pre-determined database [39]. This augmentation can occur at any stage—before, during, or after response generation—and can be iterative, improving over multiple cycles until the response is verified as accurate. Self-refinement strategies, on the other hand, utilize the model itself to detect and refine potential hallucinations.

### v2

@@ 5.2 Impact of Model Settings (RQ2)
Effect of Temperature Settings. The temperature setting in LLMs adjusts the randomness of the generated responses, where a lower temperature results in more predictable and deterministic outputs, and a higher temperature increases creativity and diversity in the responses (Section F.2). We varied this setting for each model between the minimum and maximum allowed values and observed the change in hallucination rate (the maximum temperature for GPT series models is limited to 2, while the open-source models can be set up to 5). All models exhibited a clear increase in hallucination rate as temperature value increases, with the effect becoming severe at maximum values. The OpenAI models, as shown in Figure 3, displayed only a slight increase in hallucination rate between temperatures 0 and 1, which then increased sharply between 1 and 2. Most notably, GPT-4 resulted in a hallucination rate (8.9%) nearly 4x lower than GPT 3.5 (31.8%) at its maximum temperature. At the highest temperature values, open-source models start to generate more hallucinated packages than valid packages. Most LLMs operate at a default temperature in the range of 0.7 to 1, however, our results indicate that a lower temperature value can reduce package hallucinations, with the optimal value varying per model. A lower temperature also yields more deterministic responses, presenting a trade-off between risk of hallucination and creativity. Therefore, selecting the appropriate temperature value is not a straightforward decision.
@@
Effect of Decoding Strategies. Next, we adjusted several decoding parameters (top-$p$, top-$k$, and min-$p$ values) to reduce the chances of a low probability token being selected as a potential package, with the intuition that lower probability tokens correspond to higher probabilities of hallucination in this context. Below is a summary of the parameters and the values we modified. We evaluated each listed value in isolation, followed by a combined evaluation of the values highlighted in bold, resulting in a total of 10 tests. Note that top-$k$ and min-$p$ were only tested for DeepSeek and CodeLlama, as those values are not modifiable through the OpenAI API.
@@
Varying the decoding values induced a slight increase (1.16% on average) in hallucination rate for all four models across all values tested. As we will expand on in RQ3, package hallucinations are often persistently repeated across many iterations.
This suggests that greedy decoding strategies, which prioritize the most probable tokens (i.e., the most probable token is always selected), would still generate fictitious packages. This differs from other types of hallucinations, which generally occur when low-probability tokens are sampled. This persistent nature of package hallucinations highlights the inherent complexity of the problem.
Figure 4: Hallucination rates of recent vs. all-time data sets.
Recency of Subject Matter. As described in Section 4.1, we separated our coding prompts into two temporal datasets to evaluate if hallucination rate was correlated with topics/packages that emerged after the model was trained. A lower difference between the rates of recent and all-time prompts indicates better performance in handling questions that fall outside the model’s pre-training data and therefore a more generalizable model. The models we tested were shown to be more likely to generate a package hallucination when responding to prompts that deal with more recent topics. This difference resulted in a 10% higher hallucination rate on average for older data vs. more recent data.
Overall, all 16 Python models we evaluated demonstrated a higher hallucination rate when being prompted about questions or packages that were popular within the past year (Figure 4). These higher rates are at least partly due to inherent limitations and training costs of modern LLMs. As noted in the OpenAI GPT-4 technical report [1], LLMs cannot update themselves with new information post-release and have no knowledge of the world beyond their training data cut-off date. Although fine-tuning can enhance specific tasks, it does not generally improve the model’s overall knowledge of the world. The massive cost of training LLMs from scratch, evidenced by the 1,400,000 GPU hours (220 years) required to train the 12 CodeLlama models, makes continuously updating pre-training data impractically expensive [66]. This cost, along with steadily increasing model sizes and training times, poses a significant barrier to reducing package hallucinations for advanced coding prompts and packages.
RQ2 Summary: Lower temperatures result in the lowest hallucination rate while hallucination rates increase dramatically with temperature values larger than 1. Altering decoding and sampling parameters in the model does not improve hallucination rates. More recent prompting topics lead to a 10% increase in hallucination rates.
@@ 5.3 Model Behaviors (RQ3)
Frequency of Repeated Hallucinations. To determine whether hallucinations are random error or repeatable phenomena, this test focuses on the persistence of hallucinations within a model. We randomly sampled 500 prompts that generated package hallucinations during our initial testing, then repeated those queries 10 times per prompt. Of those 10 queries we recorded how many times the original hallucinated package was regenerated. Our analysis reveals an unexpected dichotomy when repeatedly querying a model with the same prompt that generated a hallucination: 43% of hallucinated packages were repeated in all 10 queries, while 39% did not repeat at all across the 10 queries. This is indicated in Figure 5, which shows prominent spikes at zero repetitions and at 10 repetitions respectively for all models. Further, 58% of the time, a hallucinated package is repeated more than once in 10 iterations, which shows that a majority of hallucinations are not simply random errors, but a repeatable phenomenon that persists across multiple iterations. This is significant because a persistent hallucination is more valuable for malicious actors looking to exploit this vulnerability and makes the hallucination attack vector a more viable threat.
Figure 5: Frequency of an identical hallucinated package name generated from the same prompt across 10 trials.
Verbose Models vs. Conservative Models. LLMs operate with inherent randomness and uncertainty. This behavior enables novel and creative output, a desired feature for many NLP tasks but less welcome for code generation, which requires a high degree of accuracy and must adhere to rigid syntax. We define a verbose model as one that operates with higher degree of uncertainty and randomness by generating a greater number of distinct package names while a conservative model generates a lesser number of distinct packages, generally using only the most popular and well-known packages. To this end, we investigated whether verbose models correspond to higher rate of package hallucinations. Our results show a correlation between the hallucination rate and number of unique packages that were recommended during this experiment (i.e. a more verbose model was associated with a higher hallucination rate). In light of these findings, as shown in Figure 6, it is reasonable to suggest that coding models, or models performing coding tasks, should operate in a more conservative manner when generating packages to answer a coding prompt.
@@
Curiously, 3 of the 4 models (GPT 4 Turbo, GPT 3.5, and DeepSeek) proved to be highly adept at detecting their own hallucinations (see Figure 7) with detection accuracy above 75%. The precision and recall values for this test, which also averaged above 75%, are presented in Table 2. This phenomenon implies that each model’s specific error patterns are detectable by the same mechanisms that generate them, suggesting an inherent self-regulatory capability. The indication that these models have an implicit understanding of their own generative patterns which could be leveraged for self-improvement is an important finding for developing mitigation strategies.
Table 2: Performance of hallucination detection tests.
CodeLlama displays unique and interesting behavior during both the tests, as it has an overwhelming propensity to label most packages as valid, resulting in a lower accuracy for hallucinated packages.
RQ3 Summary: Package hallucinations are often persistently generated. Models that generate fewer packages when prompted are correlated with a reduced hallucination rate. Several models were able to accurately detect their own hallucinations.
@@ 5.4 Characteristics of Hallucinations (RQ4)
Occurrence of the Same Package Hallucination Across Different Models. To analyze the possibility of the same hallucinated packages being generated across different models, we measured how many models generated the same package name given a confirmed package hallucination. Figure 8 shows a clear pattern where a vast majority (81%) of distinctly generated package names were generated by only one model. In other words, specific package names were usually unique to a single model, where only the most common packages were generated by more than one model. The two populations diverge as the number of models increases, with the number of hallucinated packages decreasing nearly exponentially and the distribution of valid packages becoming more uniform. The finding that valid packages are less dependent on the specific model used for generation is attributable to their frequent appearance in training data and applicability to universal coding problems, leading to their widespread use in a broad range of prompts.
@@
Combining the insights gained during the previously discussed persistence analysis (Figure 5) leads to a key observation. As previously shown, hallucinations are often persistent (58% are repeated within 10 iterations) within the same model but are not often repeated between models, as 81% of hallucinated packages are generated by only one model. This further reinforces the evidence that while hallucinations are a common phenomenon across various models, the exact nature of these hallucinations is generally model-specific. This behavior is particularly surprising given that our testing includes multiple models from the same family (i.e. 3 GPT models, 4 CodeLlama models, and 3 DeepSeek models). These models presumably use the same training data for each version, yet each model generates unique hallucinations not found in other models.
Figure 8: Number of models in which each unique package (valid & hallucinated) appeared, with the y-axis on a log scale.
@@
Semantic Similarity Between Hallucinated and popular packages. In order to analyze the semantic similarity between hallucinated and popular real/valid packages, we measured the average Levenshtein distance of a package to its nearest neighbor (i.e., the closest valid package). Levenshtein distance is a measure of how many insertions, deletions, and substitutions are required for two strings to match [36]. If the distribution of Levenshtein distances is skewed heavily right, with a peak at or near 0, this would indicate that most hallucinations are very similar to valid package names. In that case, attackers could infer a hallucination target based on more traditional package confusion methods (e.g., typosquatting) rather than analyzing a large volume of model output over time to detect persistent hallucinations that could be used as vessels for malicious code. A higher distance reflects that package hallucinations are more random in nature and difficult to predict, rather than the result of minor grammatical errors.
@@
The results of the Levenshtein distance test, as seen in Figure 9, suggests that most package hallucinations are not simple off-by-one errors. An off-by-one error in our case is defined as a difference between the hallucinated package and its closest match of 1-2 numbers, letters, or punctuation marks, with only 13.4% (10,263 of 76,489) belonging to this category. Another 37.9% (29,025 of 76,489) of packages registered a score between 3-5, which would indicate two words with a common root word or concept that still differ significantly. Notably, 48.6% (37,207 of 76,489) of hallucinations scored 6 or greater, with 20.2% (15,457 of 76,489) of those scoring 10 or higher, indicating two strings that are very different and likely do not share any common theme.
@@
The presence of such a large proportion of high Levenshtein values suggests that the majority of hallucinations are not merely trivial typographical errors but are instead substantively different from existing package names. The observed results provide further evidence that the root cause of hallucinations is likely to be more complex than minor string manipulation, pointing to deeper issues in the model’s generative processes that govern the creation of package names. The long right tail of the distribution in Figure 9 indicates a wide variety of hallucinations spread across a broad range, revealing a diversity in types of errors and reinforcing that the generation of hallucinations is a complex issue not limited to simple character substitutions, additions, and deletions.
@@
Effect of Language Confusion. Another behavior observed during the main experiment is the tendency to confuse programming languages while generating package output (i.e. the model is asked to provide Python packages, but instead provides JavaScript packages). To validate the potential existence of this behavior in code generating LLMs, we obtained master lists of packages (using libraries.io [72]) from the nine most popular open-source repositories and compared our list of hallucinated package names generated during Python testing to the respective master lists of valid packages from other programming languages. Any intersection between the two lists indicates a cross-language hallucination. Overall, only JavaScript is a significant source of cross-language hallucinations, as 8.7% (6,705/76,489) of hallucinated Python packages are valid JavaScript packages. All other languages contributed negligible hallucinations, combining for only 0.8% (663 of 76,489) across eight other open-source repositories, including R, Rust, Ruby, PHP, Swift and .NET (see Table 3 for complete results).
Table 3: Confusion by programming language repository.
RQ4 Summary: Most unique hallucinated package names were generated by a single model and thus only appeared in 1 out of 16 tests. Most hallucinated package names were not semantically similar to a valid Python package as measured by Levenshtein distance. Deleted packages were a negligible source of package hallucinations while JavaScript was the only significant source of cross-language hallucinations.
Figure 9: Levenshtein distance of hallucinated packages to nearest valid package.
@@ 6 Mitigation
Motivated by the finding of systemic package hallucination in all the tested models, we investigate techniques to mitigate the occurrence of such hallucinations (thus addressing RQ5).
@@ 6.1 Mitigation Strategies
A seemingly obvious mitigation strategy when detecting package hallucinations might involve comparing a master list of valid packages to the model’s code output. It is imperative to understand why this approach is merely a superficial solution rather than addressing the root cause. The fundamental vulnerability inherent in code generation by LLMs regarding package hallucinations is not solely that they produce non-existent packages, but that they generate package names that were not present in their training data. Our study emphasizes the detection of hallucinated packages as indicators of an underlying vulnerability. The genuine threat posed by package hallucinations emerges when such a hallucinated package is officially published and contains malicious code. The mere act of publishing does not render a previously hallucinated package benign; on the contrary, it becomes an active threat seeking to exploit a security vulnerability. Consequently, simply cross-referencing generated packages with a master list would erroneously validate a malicious package after publication. Dedicated mitigation techniques are essential to prevent the generation of hallucinated packages altogether, thereby addressing the root cause of this latent vulnerability.
Implementing mitigation strategies specific to code generation LLMs is an unexplored topic at the time of writing, therefore we rely on general hallucination mitigation strategies proposed for standard NLP tasks that can be applied to code generation. These strategies can generally be grouped into two broad categories: prompt engineering and model development [73]. Prompt engineering includes methods such as Retrieval Augmented Generation (RAG), self-refinement, and prompt tuning. RAG approaches involve enriching the original prompt with additional information gathered from an external source, such as the web or a pre-determined database [37]. This augmentation can occur at any stage—before, during, or after response generation—and can be iterative, improving over multiple cycles until the response is verified as accurate. Self-refinement strategies, on the other hand, utilize the model itself to detect and refine potential hallucinations.